a student opens a mathematics book to two facing pages. the product of the page numbers is 272. find the page numbers.
Answer by edjones(8007)

Let the pages be x and x+1
x(x+1)=272
x^2+x-272=0
(x+17)(x-16)=0
x=16
x+1=17
.
Ed